[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Dec 9, 2015

Dr. Craig A. Mattocks craig.mattocks at noaa.gov
Wed Dec 9 13:00:33 UTC 2015


Here is today's update on NHC's computer operations. 

Craig
-----

TAFB:

1. Marc Laurent replaced the handset on the telephone at the TAFB Atlantic desk, which fixed the disconnection problems.

2. Linux SysAdmin Martin Graham set up the palmar, atlmar, etc. accounts on the new standing N-AWIPS workstation (derecho), but the forecasters are unable to run their N-AWIPS applications. Matt Sardi will investigate.

HSU:

1. Ongoing: P-Surge test cases will be run once per day at about 14 UTC (9 am local time) all week to test the AWIPS software at WFOs and continue collaboration testing.

Other Issues:

GOES-E Sounder Update (from today's SDM log):

1. 1936Z - ESPC emailed a update on the GOES-E sounder, stating:
"The GOES-13 Sounder instrument anomaly investigation revealed that filter wheel motor has come to a complete or near complete stop. Power cycle of the motor windings has been executed for past two weeks with no observable improvement. GOES-13 (GOES-East) Sounder IR data outage will continue indefinitely. Engineers will continue to perform Sounder recovery activities. This is the final update for GOES-13 (GOES-East) Sounder IR Data Outage."
 



More information about the Ncep.nhc.nco_contacts mailing list